OpenStudy (anonymous):

find two points that lie on the line and find the slope of the line? -3x + 5y = 30

OpenStudy (anonymous):

solve for y 5y=30+3x y=6+(3/5)x The slope of the line is 3/5

OpenStudy (anonymous):

the points are (0, 6) and (-2, -1)

OpenStudy (anonymous):

to fin the points just substitute any value for x and solve for y

OpenStudy (anonymous):

thank you...can i ask a few more?

OpenStudy (anonymous):

of course

OpenStudy (anonymous):

same question as the other two. -7x - 2y = 56

OpenStudy (anonymous):

-56-7x=2y -28-2/7x=y The slope is -2/7 and the points are (0, -28) and (7, -30)

OpenStudy (anonymous):

thank you! -12x + 3y = -72

OpenStudy (anonymous):

3y=12x-72 y=4x-24 The slope is 4 and points are (0, -24) and (6, 0)

OpenStudy (anonymous):

thank you, last one 13x - 5y = -130

OpenStudy (anonymous):

13x+130=5y 13/5x+26=y The slope is 13/5 and points are (0, 26) and (-5, 13)

OpenStudy (anonymous):

thank you so so so much!!!!